# A Roadmap to Guide the Integration of LLMs in Hierarchical Planning

### v1

@@ Introduction
Hierarchical Planning (HP) is a subfield within Automated Planning (AP) comprising planning methods that incorporate hierarchical knowledge. This hierarchical knowledge can be leveraged to speed up planning and, also, to integrate human expert problem-solving knowledge. While Large Language Models (LLMs) are being gradually integrated in various AI domains, including AP, their application to HP remains underexplored, with only a few studies tangentially addressing this topic (Yang, Zhang, and Hou 2024; Dai et al. 2024; Tse 2024; Song et al. 2023). Our contribution within this work, therefore, is a roadmap to fill this gap, exploring the potential of LLMs for HP.
We have analyzed existing literature on AP with LLMs, as well as current reviews (Pallagani et al. 2024; Huang et al. 2024; Valmeekam et al. 2022). Observing that most methods in AP that are similarly applicable to HP, we have classified these methods, building a taxonomy to bring HP closer to the existing LLM integration techniques. Our taxonomy classifies the integration methods according to two different dimensions: the Planning Process Role (in which part of the HP life cycle is the LLM applied – i.e. problem definition, plan elaboration or post-processing) and the LLM Improvement Strategy (which LLM-based approach are used to improve the LLM performance – i.e. giving extra knowledge or making multiple calls). Given that HP is a subset of AP, this classification is broadly applicable to AP as well.
To facilitate evaluation and comparisons among methods, we propose a standardized dataset and benchmarking framework based on the 2023 International Planning Competition (IPC-2023) HTN tracks, the most recent competition for HP solvers. Specifically, we suggest using the total-order track of the IPC-2023 dataset as a benchmark dataset. [^1] Website of the IPC-2023: https://ipc2023-htn.github.io As a baseline, we implement and evaluate a basic LLM Planner (direct planning using an LLM without any improvement strategies), the simplest method of our taxonomy, on this dataset. We use Llama-3.1-Nemotron-70B-Instruct (Wang et al. 2024), one of the highest-performing LLMs available, and the model that we plan to use in subsequent experiments. We also provide the results from PandaDealer-agile-lama (Olz, Höller, and Bercher 2023), the IPC-2023 Total-Order Satisficing track winner and state-of-the-art HP solver.
@@
An HP Problem includes the same elements as an AP problem (Actions, Initial State and Goal), along with the high-level actions (Tasks) that represents the hierarchical information about the environment. Each element can be generated using an LLM through two main approaches: Translation, when all necessary information is explicitly and previously provided and we only want the LLM to restructure the provided information into a target format (Liu et al. 2023); and Generation, when the information is partially or implicitly provided, so the LLM is inferring or assuming missing information based on reasoning (Gestrin, Kuhlmann, and Seipp 2024).
@@
  • LLM Planner. In this basic setup, the LLM itself functions as the planner (Silver et al. 2022). While LLM improvement strategies can be applied, there is not external planner or explicit search process here, so the LLM the model is responsible for the entire planning process.
  • Graph Search. The LLM is embedded in the planner, which is done within an explicit search algorithm, where the LLM can perform one or more roles, like Node Expansion (generating the next possible actions), Selection (choosing the next action), Heuristic provision (scoring states), Model Elicitation, Backtracking, Aggregation and Pruning. Some remarkable architectures are RAP (Hao et al. 2023), integrating an LLM into a MCTS Planner; GoT (Besta et al. 2024), implementing an LLM in most of the mentioned roles to solve reasoning problems; and SayCan (Ahn et al. 2022), which utilizes an LLM as a probabilistic relevance (heuristic) scorer.
  • Planning Guidance. Here, the LLM is external to the planner, but assists it by providing assistance. This guidance can be an initial plan that the planner can refine (Valmeekam et al. 2023) or environment preferences to narrow the search space (Sharan et al. 2023).
@@
Once a plan is developed, an LLM can be used to refine it, by Translating the plan into another data structure or language, typically an executable format (e.g. for a robot to run it) or natural language (Liu et al. 2023). Or Explaining the Plan, where de LLM has to generate more detailed information, commonly in natural language, based on the provided plan (Simon and Muise 2022).
@@
  • Previous. This approach involves providing extra information to the LLM before its execution. Fine-tuning is the traditional method in the field of deep learning, which involves adjusting the model’s internal weights with supplementary data (Pallagani et al. 2022). Alternatively, knowledge can be directly provided through the model’s prompt, which is a more flexible and accessible option than fine-tuning, as it does not require a training process. Providing extra information is achieved through input-output examples (shots) that enhance generated responses (Song et al. 2023). If the shots provide useful environmental information (e.g., domain information) this is called in-context prompting; if they only aim to elicit an specific output structure, it is referred to as out-of-context prompting. Furthermore, we can also provide knowledge about how to reason, about the problem itself. Chain of Thoughts (CoT) (Wei et al. 2022) is the best example of this, a different prompting strategy that encourages the LLM to generate a reasoning process before providing the final answer. This is typically achieved by using few-shot prompting with reasoning examples, though reasoning can also be elicited without examples, known as Zero-shot CoT (Kojima et al. 2022).
  • Feedback. Unlike the previous paradigm, here the extra information is provided during the task solving process: rather than simply accepting the output of an LLM execution, we can iteratively improve the result of the LLM by providing it feedback based on its previous outputs. This feedback can come from various sources (humans, the environment, an external module, another LLM, etc.). Reinforcement Learning is the classical method of learning from feedback in machine learning, where the knowledge is applied by modifying the model’s internal weights, based on a reward score (Yao et al. 2020). Alternatively, feedback can be used to dynamically adjust and improve the prompt (Prompt Correction), where prominent examples in the literature are Self-Refine (Madaan et al. 2024) and Reflexion (Shinn et al. 2024). Finally, the model’s learning can also be represented through an external Memory Module, being Voyager a notable architecture that uses this approach (Wang et al. 2023).
@@
  • Decomposition. A problem can be sequentially decomposed: each LLM call’s output generates an intermediate step that is utilized for the next call’s input. For instance, to generate a plan, we could iteratively ask the LLM to only generate the next action of a partial plan, than attempting to generate the entire plan in a single call (Huang et al. 2022). Otherwise, the problem can be decomposed in parallel: each LLM call can returns a list simpler sub-problems which can be independently solved (i.e. the classical Divide and Conquer strategy). This is illustratively used in Generative Agents (Park et al. 2023).
  • Revision. We can make the multiple calls to sequentially refine the LLM’s output: we ask the LLM to initially give us a complete but simple (general) solution that iteratively becomes more detailed through the calls (Liu et al. 2024). Otherwise, we can do a parallel approach by asking the LLM to entirely solve the problem multiple times and, then, combine the different results into a final output. Self-consistency (Wang et al. 2022) is a prominent approach, which combines the several outputs by selecting the most common response, but various criteria can be used to choose the final output.
@@
Hierarchical Task Networks (HTN) Planning (Nau et al. 2003) is the most widely used and studied approach within the HP field, making it a suitable reference point for experimenting with and evaluating various LLM integration methods. We propose using the total-order track from IPC-2023, which contains 22 domains, each with dozens of problems. A detailed breakdown of this information, along with a summary of the experimental results, is provided in Table 3.
To assess and compare the performance of future implementations, we establish two reference points. First, we consider the winner of the IPC-2023 Total-Order Satisficing track, PandaDealer-agile-lama (Olz, Höller, and Bercher 2023) as an upper bound, as it represents the current state of the art in HP solvers. Second, we include as baseline a basic LLM Planner (i.e. using an LLM directly to plan without providing any LLM Improvement Strategy) representing a lower bound for LLM Integration. This approach, being the simplest method in our taxonomy, serves as a foundational point of comparison and a starting point for the roadmap.
@@
We utilize Llama-3.1-Nemotron-70B-Instruct, one of the highest-performing LLMs available (Wang et al. 2024), as the LLM Planner. We utilize the official IPC verifier to score the generated plans along the different proposed metrics. The execution and validation source codes, as well as the generated plans, and obtained results are allocated on GitHub. URL: https://github.com/Corkiray/HTN-LLM.

### v2

@@ Introduction
Hierarchical Planning (HP) is a subfield within Automated Planning (AP) comprising planning methods that incorporate hierarchical knowledge. This hierarchical knowledge can be leveraged to speed up planning and, also, to integrate human expert problem-solving knowledge. While Large Language Models (LLMs) are being gradually integrated in various AI domains, including AP, their application to HP remains underexplored, with only a few studies tangentially addressing this topic (yang2024oceanplan; dai2024optimal; tseimproving; song2023llm). Our contribution within this work, therefore, is a roadmap to fill this gap, exploring the potential of LLMs for HP.
We have analyzed existing literature on AP with LLMs, as well as current reviews (pallagani2024prospects; huang2024understanding; valmeekam2022large). Observing that most methods in AP that are similarly applicable to HP, we have classified these methods, building a taxonomy to bring HP closer to the existing LLM integration techniques. Our taxonomy classifies the integration methods according to two different dimensions: the Planning Process Role (in which part of the HP life cycle is the LLM applied – i.e. problem definition, plan elaboration or post-processing) and the LLM Improvement Strategy (which LLM-based approach are used to improve the LLM performance – i.e. giving extra knowledge or making multiple calls). Given that HP is a subset of AP, this classification is broadly applicable to AP as well.
To facilitate evaluation and comparisons among methods, we propose a standardized dataset and benchmarking framework based on the 2023 International Planning Competition (IPC-2023) HTN tracks, the most recent competition for HP solvers. Specifically, we suggest using the total-order track of the IPC-2023 dataset as a benchmark dataset. [^1] Website of the IPC-2023: https://ipc2023-htn.github.io As a baseline, we implement and evaluate a basic LLM Planner (direct planning using an LLM without any improvement strategies), the simplest method of our taxonomy, on this dataset. We use Llama-3.1-Nemotron-70B-Instruct (wang2024helpsteer2preferencecomplementingratingspreferences), one of the highest-performing LLMs available, and the model that we plan to use in subsequent experiments. We also provide the results from PandaDealer-agile-lama (olz2023pandadealer), the IPC-2023 Total-Order Satisficing track winner and state-of-the-art HP solver.
@@
An HP Problem includes the same elements as an AP problem (Actions, Initial State and Goal), along with the high-level actions (Tasks) that represents the hierarchical information about the environment. Each element can be generated using an LLM through two main approaches: Translation, when all necessary information is explicitly and previously provided and we only want the LLM to restructure the provided information into a target format (liu2023llm); and Generation, when the information is partially or implicitly provided, so the LLM is inferring or assuming missing information based on reasoning (gestrin2024nl2plan).
@@
  • LLM Planner. In this basic setup, the LLM itself functions as the planner (silver2022pddl). While LLM improvement strategies can be applied, there is not external planner or explicit search process here, so the LLM the model is responsible for the entire planning process.
  • Graph Search. The LLM is embedded in the planner, which is done within an explicit search algorithm, where the LLM can perform one or more roles, like Node Expansion (generating the next possible actions), Selection (choosing the next action), Heuristic provision (scoring states), Model Elicitation, Backtracking, Aggregation and Pruning. Some remarkable architectures are RAP (hao2023reasoning), integrating an LLM into a MCTS Planner; GoT (besta2024graph), implementing an LLM in most of the mentioned roles to solve reasoning problems; and SayCan (ahn2022can), which utilizes an LLM as a probabilistic relevance (heuristic) scorer.
  • Planning Guidance. Here, the LLM is external to the planner, but assists it by providing assistance. This guidance can be an initial plan that the planner can refine (valmeekam2023planning) or environment preferences to narrow the search space (sharan2023llm).
@@
Once a plan is developed, an LLM can be used to refine it, by Translating the plan into another data structure or language, typically an executable format (e.g. for a robot to run it) or natural language (liu2023llm). Or Explaining the Plan, where de LLM has to generate more detailed information, commonly in natural language, based on the provided plan (simon2022tattletale).
@@
  • Previous. This approach involves providing extra information to the LLM before its execution. Fine-tuning is the traditional method in the field of deep learning, which involves adjusting the model’s internal weights with supplementary data (Pallagani_Muppasani_Murugesan_Rossi_Horesh_Srivastava_Fabiano_Loreggia_2022). Alternatively, knowledge can be directly provided through the model’s prompt, which is a more flexible and accessible option than fine-tuning, as it does not require a training process. Providing extra information is achieved through input-output examples (shots) that enhance generated responses (song2023llm). If the shots provide useful environmental information (e.g., domain information) this is called in-context prompting; if they only aim to elicit an specific output structure, it is referred to as out-of-context prompting. Furthermore, we can also provide knowledge about how to reason, about the problem itself. Chain of Thoughts (CoT) (Wei_Wang_Schuurmans_Bosma_Ichter_Xia_Chi_Le_Zhou_2023) is the best example of this, a different prompting strategy that encourages the LLM to generate a reasoning process before providing the final answer. This is typically achieved by using few-shot prompting with reasoning examples, though reasoning can also be elicited without examples, known as Zero-shot CoT (kojima2022large).
  • Feedback. Unlike the previous paradigm, here the extra information is provided during the task solving process: rather than simply accepting the output of an LLM execution, we can iteratively improve the result of the LLM by providing it feedback based on its previous outputs. This feedback can come from various sources (humans, the environment, an external module, another LLM, etc.). Reinforcement Learning is the classical method of learning from feedback in machine learning, where the knowledge is applied by modifying the model’s internal weights, based on a reward score (yao2020keep). Alternatively, feedback can be used to dynamically adjust and improve the prompt (Prompt Correction), where prominent examples in the literature are Self-Refine (madaan2024self) and Reflexion (Shinn_Cassano_Berman_Gopinath_Narasimhan_Yao_2023). Finally, the model’s learning can also be represented through an external Memory Module, being Voyager a notable architecture that uses this approach (Wang_Xie_Jiang_Mandlekar_Xiao_Zhu_Fan_Anandkumar_2023).
@@
  • Decomposition. A problem can be sequentially decomposed: each LLM call’s output generates an intermediate step that is utilized for the next call’s input. For instance, to generate a plan, we could iteratively ask the LLM to only generate the next action of a partial plan, than attempting to generate the entire plan in a single call (Huang_Abbeel_Pathak_Mordatch_2022). Otherwise, the problem can be decomposed in parallel: each LLM call can returns a list simpler sub-problems which can be independently solved (i.e. the classical Divide and Conquer strategy). This is illustratively used in Generative Agents (park2023generative).
  • Revision. We can make the multiple calls to sequentially refine the LLM’s output: we ask the LLM to initially give us a complete but simple (general) solution that iteratively becomes more detailed through the calls (liu2024delta). Otherwise, we can do a parallel approach by asking the LLM to entirely solve the problem multiple times and, then, combine the different results into a final output. Self-consistency (wang2022self) is a prominent approach, which combines the several outputs by selecting the most common response, but various criteria can be used to choose the final output.
@@
Hierarchical Task Networks (HTN) Planning (nau2003shop2) is the most widely used and studied approach within the HP field, making it a suitable reference point for experimenting with and evaluating various LLM integration methods. We propose using the total-order track from IPC-2023, which contains 22 domains, each with dozens of problems. A detailed breakdown of this information, along with a summary of the experimental results, is provided in Table 3.
To assess and compare the performance of future implementations, we establish two reference points. First, we consider the winner of the IPC-2023 Total-Order Satisficing track, PandaDealer-agile-lama (olz2023pandadealer) as an upper bound, as it represents the current state of the art in HP solvers. Second, we include as baseline a basic LLM Planner (i.e. using an LLM directly to plan without providing any LLM Improvement Strategy) representing a lower bound for LLM Integration. This approach, being the simplest method in our taxonomy, serves as a foundational point of comparison and a starting point for the roadmap.
@@
We utilize Llama-3.1-Nemotron-70B-Instruct, one of the highest-performing LLMs available (wang2024helpsteer2preferencecomplementingratingspreferences), as the LLM Planner. We utilize the official IPC verifier to score the generated plans along the different proposed metrics. The execution and validation source codes, as well as the generated plans, and obtained results are allocated on GitHub. URL: https://github.com/Corkiray/HTN-LLM.
@@ Conclusion
Acknowledgment
This work has been partially funded by the Grant PID2022-142976OB-I00, funded by MICIU/AEI/ 10.13039/501100011033 and by “ERDF/EU”.